### v6

@@ 1 Introduction
Recently, the utilization of diffusion models (DMs) (Sohl-Dickstein et al., 2015; Ho et al., 2020; Ramesh et al., 2021) and large language models (LLMs) (Radford et al., 2019; Brown et al., 2020; OpenAI, 2023) has surged across various real-world applications (Ramesh et al., 2021; Saharia et al., 2023; OpenAI, 2023), enabling the generation of content that rivals human expertise (Ho et al., 2020; OpenAI, 2023). For instance, ChatGPT, a milestone LLM developed by OpenAI, has become a ubiquitous productivity tool worldwide, offering invaluable assistance across diverse domains—from serving as a virtual assistant to generating code segments for computer engineers (Brown et al., 2020; OpenAI, 2023). Moreover, multimodal models built upon DMs and LLMs have achieved significant advancements in transforming content from one modality to another, particularly in bridging language and vision (Ramesh et al., 2021, 2022; Xu, 2022; Avrahami et al., 2022; Harvey et al., 2022). This can be exemplified by Stable Diffusion (Rombach et al., 2022), which converts textual descriptions into visual art.
@@
While these models offer substantial social benefits, their malicious exploitation (Carlini et al., 2023a, 2021; Tamkin et al., 2021) has raised significant concerns about their trustworthiness. DMs have been criticized for exacerbating societal divisions, as the images they generate may revive harmful stereotypes or manipulate public opinion. For example, in April 2023, an organization misused DMs to generate misleading information for specific agendas. Similarly, LLMs have been implicated in serious issues, including contributing to suicide cases, fabricating legal cases, and leaking users’ chat histories. This troubling trend is expected to escalate, with an official academic institution suggesting that within a few years, 90% of online content could be generated by these models. A report from the World Economic Forum predicts that such content will completely reshape public perception in the near future. In response, the trustworthiness landscape of DMs and LLMs is evolving rapidly, with numerous initiatives underway. However, a significant gap remains in systematically organizing and critically reviewing these efforts. To fill this gap, as shown in Figure 1, we embark on a systematic trustworthiness investigation by organizing recent advances around four fundamental dimensions: privacy, security, fairness, and responsibility:
@@
Distinct features of this survey. The latest surveys that are most closely related to this survey include (Liu et al., 2023d; Huang et al., 2023c; Cao et al., 2024). Cao et al. (2024) focused on the trustworthiness of DMs, while Liu et al. (2023d) evaluated LLM alignment with human behavior. Huang et al. (2023c) scrutinized existing research from a benchmarking perspective. In contrast, our survey extends beyond the scope of these papers in two significant ways. First, it expands the horizon of current surveys by amalgamating insights on both DMs and LLMs, aligning with the prevailing trend of integrating these as multimodal models. By comparing the trustworthiness of LLMs and DMs, we aim to deepen the understanding of the distinct characteristics inherent in each modality, fostering interdisciplinary dialogue and encouraging the exchange of methodologies and theoretical insights. Second, this survey goes beyond engineering-focused metrics and standardized benchmarks to provide a comprehensive understanding of the development and evolution of trustworthiness in DMs and LLMs, advocating for flexible evaluation procedures that account for regional and temporal variations. In summary, this survey yields four key benefits:
@@ 2 A Glimpse of State-of-the-art Generative Models
DMs. At the core of DMs lies the diffusion process, inspired by non-equilibrium thermodynamics (Sohl-Dickstein et al., 2015; Song et al., 2021), which gradually transforms a simple distribution, typically Gaussian noise, into a complex distribution. Formally, given a natural sample $x_{0}$, the transition between two consecutive diffusion steps is defined as:
@@
This iterative adjustment simplifies the learning process compared to generating an image in a single step, allowing for richer textures and more intricate details in the final outputs (Saharia et al., 2022c). The vanilla diffusion model (Sohl-Dickstein et al., 2015) is designed for unconditional image generation tasks. For conditional tasks (Ramesh et al., 2021), additional supervision signals $y$ guide the models to produce images that correspond $y$, enhancing the adaptability and versatility of the models across various tasks (Harvey et al., 2022; Saharia et al., 2022a). This requires an encoder $\tau_{w}$ to map $y$ to a latent vector, which is incorporated into the diffusion and generation processes, i.e., modifying $\epsilon_{\theta}(x_{t},t)$ to $\epsilon_{\theta}(x_{t},\tau_{w}(y),t)$. Among these supervision signals, labels and textual descriptions (Ramesh et al., 2021) are particularly significant, enabling DMs to generate images that align with the provided descriptions. For tasks like image coloring, supervision may also come in the form of images, where the model transforms a gray-scale source image into its colored counterpart (Saharia et al., 2022a).
@@ 3 Privacy
The extensive number of parameters in DMs and LLMs enables them to learn from vast amounts of training data. However, this over-parameterization can unintentionally create shortcuts that allow models to achieve high performance by merely memorizing training samples, leading to privacy leakage. Example 1 illustrates a real-world incident of privacy leakage involving a DM, inciting significant criticism from Getty Images, the owner of some of the implicated training data. This underscores the urgent need to address privacy issues when deploying models, particularly in sensitive sectors such as healthcare. We review recent developments concerning the privacy issues of DMs and LLMs.
@@ 3.1 Overview
A model is considered privacy-preserving if it safeguards information about its training data throughout its entire lifecycle, in a way that no feasible methods exist to derive such information. This definition aligns with the general principles of data protection as outlined in legal frameworks across various countries [^2] https://www.dlapiperdataprotection.com/. Owing to the distinct differences between training and inference stages, we examine data leakage concerns separately for each phase. During the training stage, we focus on federated learning and split learning, while in Section 3.4, we specifically address membership inference attacks targeting DMs and LLMs, which seek to determine the membership status of particular data.
@@ 3.2 Data Leakage during Training Stage
Two paradigms exist for training deep neural networks: centralized training and distributed training with multi-party participation (Li et al., 2020a). Centralized training offers better data protection, especially when robust access control measures are in place. Conversely, the latter can heighten the risk of privacy leakage due to the potential involvement of untrusted parties. Federated learning and split learning are two key collaborative learning frameworks. Unless stated otherwise, we will not differentiate between DMs and LLMs in this context, as both federated learning and split learning are frameworks that can be applied to either model type.
@@ 3.2.1 Federated Learning
In federated learning, a server distributes a globally shared model to clients, who then perform local computations on their individual raw data to generate gradients. These gradients are sent back to the server, which aggregates them to update the global model. While federated learning appears to protect privacy by forgoing the need to share raw data, gradient leakage attacks have shown that the shared gradients can still be sufficient to reconstruct clients’ raw data.
@@
Gradient Leakage Attacks. Gradient leakage attacks (Zhu et al., 2019; Zhao et al., 2020) exploit gradient matching techniques to recover clients’ data. Wherein, attackers align the uploaded gradients with the gradients generated by dummy data, enabling high-fidelity data reconstruction. Subsequent research (Geiping et al., 2020; Yin et al., 2021; Wei et al., 2020) has improved the efficacy of these attacks by utilizing cosine similarity loss functions, adding regularization terms, and refining initialization methods, among others. Interestingly, optimized dummy images often resemble random noise, which hinders attack performance (Jeon et al., 2021). To address this, Jeon et al. (2021) and Yue et al. (2023) optimized generative models’ latent vectors to generate clearer images. Deng et al. (2021) found that a hybrid loss function combining $L_{1}$ and $L_{2}$ norms can enhance text reconstruction results. Balunovic et al. (2022) proposed a unified Bayesian framework for gradient leakage attacks, noting that larger models are more susceptible to privacy leakage. Besides, DMs and LLMs tend to be more susceptible to gradient leakage attacks because they directly model the training data. For example, Gupta et al. (2022) leveraged the generative capabilities of these models to generate candidate data and determined the optimal reconstruction by comparing the similarity between candidate gradients and the uploaded gradients. As model performance improves, so does the quality of reconstruction.
@@
Gradient Leakage Defenses. Encryption and perturbation are the primary techniques used to defend against gradient leakage attacks. The encryption methods (Zhang et al., 2020b; Wagh et al., 2020) prevent attackers from accessing individual client gradients, allowing only access to the plaintext of aggregated gradients, which forces them to reconstruct all clients’ data at once and complicates gradient matching problem. However, encryption methods can introduce a high computational overhead, limiting their usability in resource-constrained settings (Sun et al., 2021; Yue et al., 2023). Perturbation-based methods apply slight gradient modifications to confound attackers. Differential privacy (Abadi et al., 2016) injects random noise into gradients, while Top-K gradient sparsification (Zhu et al., 2019) retains only the most significant gradient elements. Gradient quantization (Yue et al., 2023) represents the gradients with lower-bit precision, reducing the amount of sensitive information. Selective pruning evaluates the private information within each gradient element and then implements gradient pruning (Sun et al., 2021; Wang et al., 2022). Another approach is to generate data that contains less private information, typically by optimizing metrics related to data privacy and model utility (Fan et al., 2022a).
@@
Remark. Although gradient leakage attacks and defenses were initially designed for discriminative models, they can also be applied to generative models. Moreover, DMs and LLMs face unique privacy challenges due to the ability to replicate patterns from training data, making them more vulnerable to these attacks. As models become more proficient, the risk of disclosing training data also increases. In light of this, one possible mitigation involves commencing training with sensitive data and transitioning to less sensitive data; however, the impact of this strategy on model performance remains uncertain. For LLMs, RAG can be employed to store privacy-sensitive information in a local database, excluding sensitive data from the training set and enhancing privacy protection. Furthermore, the prevalent practice of fine-tuning LLMs with adapters, as opposed to training them from scratch, allows for a restricted sharing of gradients in federated learning, reducing the risk of privacy breaches. Nonetheless, the effectiveness of the above several mitigations in counterbalancing the privacy risks introduced by models’ generation capabilities remains unclear.
@@ 3.2.2 Split Learning
In split learning (Vepakomma et al., 2018), a model is divided into two sub-models: the bottom network, which resides with the data clients, and the top network, located with the label server. Each iteration begins with the clients processing input data through the bottom network to produce intermediate results, which are then sent to the server for further computation using the top network. Privacy concerns arise when one party is deemed untrustworthy by the other.
Data Leakage. Pasquini et al. (2021) introduced the Feature Space Hijacking Attack (FSHA), targeting the reconstruction of client data when the label party is deemed untrustworthy. In FSHA, the adversarial label party trains a shadow bottom network to mimic the outputs of client’s bottom network, and then a decoder maps these outputs back to their corresponding data. By applying outputs from client’s bottom network to the trained decoder, client data can be revealed. Chen et al. (2024b) relaxed the requirement for FSHA, allowing the use of common public data through transfer learning instead of data similar to client data. Xu et al. (2024c) demonstrated that replacing the decoder with a diffusion model can yield better reconstruction results.
Existing defense measures, such as detection and perturbation, are found to be ineffective against FSHA (Pasquini et al., 2021). To address this, Li et al. (2022a) developed ResSFL, which trains a feature extractor resistant to inversion and initializes the bottom model with it, thus reducing the risk of privacy leakage. Maeng et al. (2024) theoretically explored the bound of the trained decoder’s ability to reconstruct the raw data by utilizing Fisher information. Luo et al. (2023) employed a regularization term to reduce the correlation between the input data and intermediate activation values, alongside pruning sensitive parameters for defense.
Label Leakage. Li et al. (2022b) showed that the sign patterns of backward gradients from the top model can lead to label leakage and suggested mitigating the risk with random gradient perturbations. However, Xiao et al. (2021) argued against the effectiveness of this method and advocated for multiple activations and label mixing to reduce the risk. Erdogan et al. (2021) and Kariyappa and Qureshi (2021) demonstrated that labels could be inferred with gradient matching techniques, and Fu et al. (2022) found that even a small amount of labeled samples is sufficient to fine-tune models, enabling direct label predictions for training data. Wan et al. (2023b) suggested the use of flipped labels to compute gradients to obfuscate attackers while training a private sub-model with true labels to compensate for performance losses resulting from flipped labels.
@@ 3.3.2 LLMs
Association attacks. This type of attack involves designing a template, which is then filled with entity names to prompt the model to predict the corresponding sensitive information. As an example, Lehman et al. (2021) studied the model’s vulnerability to exposing sensitive medical information using a template: ”[NAME] symptoms of [masked].” Huang et al. (2022) examined the risk of leaking personal information, such as email addresses and phone numbers. Kim et al. (2023b) enhanced association attacks by employing likelihood ratio scores for more accurate predictions. Recent studies have also explored how the integration of external retrieval data affects privacy leakage in RAG. These works (Huang et al., 2023b; Zeng et al., 2024; Qi et al., 2024; Yu et al., 2023a) crafted specific templates to compel LLMs to output the retrieved data. The templates generally consist of prompts specifying the attacker-desired content and a command directing the model to present the retrieved content.
Defenses. Data deduplication (Carlini et al., 2021; Jagielski et al., 2023; Carlini et al., 2023b; Kandpal et al., 2022) and differential privacy (Carlini et al., 2021; Lukas et al., 2023) are both applicable to LLMs and present remarkable effectiveness in practice. Nevertheless, it is worth noting that differential privacy does come with the privacy-utility trade-off. For further protection, Lukas et al. (2023) suggested utilizing named entity recognition as a method to filter out sensitive information present in the training sets. This sensitive-information-filtering method works well in RAG (Huang et al., 2023b). Moreover, another approach to boost privacy protection in RAG is by blending public and private data in datastore and encoder training (Huang et al., 2023b).

### v7

@@ 1 Introduction
The utilization of diffusion models (DMs) (Ho et al., 2020; Ramesh et al., 2021) and large language models (LLMs) (OpenAI, 2023) has surged across various real-world applications, enabling the generation of content that rivals human expertise. For instance GPT-4 has become a ubiquitous productivity tool worldwide, offering invaluable assistance across diverse domains—from serving as a virtual assistant to generating code segments for engineers (Brown et al., 2020; OpenAI, 2023). Moreover, multimodal models built upon DMs and LLMs have achieved significant advancements in transforming content from one modality to another, particularly in bridging language and vision (Ramesh et al., 2022; Avrahami et al., 2022; Harvey et al., 2022).
@@
While these models offer substantial social benefits, their malicious exploitation (Carlini et al., 2021, 2023a) has raised significant concerns about their trustworthiness. DMs have been criticized for exacerbating societal divisions, as the images they generate may revive harmful stereotypes or manipulate public opinion. For example, in April 2023, an organization misused DMs to generate misleading information for specific agendas. Similarly, LLMs have been implicated in serious issues, including contributing to suicide cases, fabricating legal cases, and leaking users’ chat histories. This troubling trend is expected to escalate, with an official academic institution suggesting that within a few years, 90% of online content could be generated by these models. A report from the World Economic Forum predicts that such content will completely reshape public perception in the near future. In response, the trustworthiness landscape of DMs and LLMs is evolving rapidly, with numerous initiatives underway. However, a significant gap remains in systematically organizing and critically reviewing these efforts. To fill this gap, as shown in Figure 1, we embark on a systematic trustworthiness investigation by organizing recent advances around four fundamental dimensions: privacy, security, fairness, and responsibility:
@@
Distinct features of this survey. Cao et al. (2024) reviewed the trustworthiness of DMs, while Liu et al. (2023d) evaluated LLM alignment with human behavior. Huang et al. (2023b) scrutinized existing research from a benchmarking perspective. In contrast, our survey extends beyond the scope of these papers in two significant ways. First, it expands the horizon of current surveys by amalgamating insights on both DMs and LLMs, aligning with the prevailing trend of integrating these as multimodal models. By comparing the trustworthiness of LLMs and DMs, we aim to deepen the understanding of the distinct characteristics inherent in each modality, fostering interdisciplinary dialogue and encouraging the exchange of methodologies and theoretical insights. Second, this survey goes beyond engineering-focused metrics and standardized benchmarks to provide a comprehensive understanding of the development and evolution of trustworthiness in DMs and LLMs, advocating for flexible evaluation procedures that account for regional and temporal variations. In summary, this survey yields four key benefits:
@@ 2 A Glimpse of State-of-the-art Generative Models
DMs. At the heart of DMs lies the diffusion process, inspired by non-equilibrium thermodynamics (Sohl-Dickstein et al., 2015; Song et al., 2021b), which gradually converts a simple distribution, typically Gaussian noise, into a complex one. Formally, given a natural sample $x_{0}$, the transition between two consecutive diffusion steps is defined as:
@@
This iterative adjustment simplifies the learning process compared to generating an image in a single step, allowing for richer textures and more intricate details in the final outputs (Saharia et al., 2022c). The above formulation is known as denoising diffusion probabilistic models (DDPM). However, the denoising process is time-consuming due to the large number of steps required. To improve efficiency, a variation called denoising diffusion implicit models (DDIM) (Song et al., 2021a) adopts a non-Markovian approach, enabling the model to skip certain steps. Other recent advances include consistency model (Song et al., 2023) and rectified flow (Liu et al., 2023c). The former learns to predict the clean data from noisy samples directly at each diffusion step, while the latter regularizes the denoising trajectory to make it simpler and more sampling-efficient. Some research proposed performing the diffusion and denoising processes in latent space instead of pixel space (Rombach et al., 2022). Moreover, while vanilla DMs are designed for unconditional image generation (Sohl-Dickstein et al., 2015), they can be extended to conditional tasks (Ramesh et al., 2021) with additional supervision signals $y$ to enhance their adaptability and versatility (Harvey et al., 2022; Saharia et al., 2022a). This requires an encoder $\tau_{w}$ to map $y$ to a latent vector, which is incorporated into the diffusion and denoising processes, i.e., modifying $\epsilon_{\theta}(x_{t},t)$ to $\epsilon_{\theta}(x_{t},\tau_{w}(y),t)$. Among these supervision signals, labels and textual descriptions (Ramesh et al., 2021) are particularly significant, enabling DMs to generate images that align with the provided descriptions. For tasks like image coloring, supervision may also come in the form of images, where the model transforms a gray-scale source image into its colored counterpart (Saharia et al., 2022a).
@@ 3 Privacy
The extensive number of parameters in DMs and LLMs enables them to learn from vast training data. However, this over-parameterization can unintentionally create shortcuts that allow models to achieve high performance by merely memorizing training samples, leading to privacy leakage. Example 1 illustrates a real-world incident of privacy leakage involving a DM, inciting significant criticism from Getty Images, the owner of some of the implicated training data. This underscores the urgent need to address privacy issues when deploying models.We review recent developments concerning the privacy issues of DMs and LLMs.
@@ 3.1 Overview
A model is considered privacy-preserving if it safeguards information about its training data throughout its entire lifecycle, in a way that no feasible methods exist to derive such information [^2] https://www.dlapiperdataprotection.com/. Owing to the distinct differences between training and inference stages, we examine data leakage concerns separately for each phase. During the training stage, we focus on federated learning and split learning, while in Section 3.4, we specifically address membership inference attacks targeting DMs and LLMs, which seek to determine the membership status of particular data.
@@ 3.2 Data Leakage during Training Stage
Two paradigms exist for training deep neural networks: centralized training and distributed training with multi-party participation (Li et al., 2020a). Centralized training provides better data protection, particularly with strong access control. Conversely, the latter can heighten the risk of privacy leakage due to the involvement of untrusted parties. Federated learning and split learning are two key collaborative frameworks. Unless specified, we do not distinguish between DMs and LLMs in this context, as both frameworks can be applied to either model type.
@@ 3.2.1 Federated Learning
In federated learning, a server distributes a globally shared model to clients, who perform local computations on their data to generate gradients. These gradients are sent back to be aggregated to update the global model. While federated learning avoids sharing raw data, gradient leakage attacks can still reconstruct clients’ data from shared gradients.
@@
Gradient Leakage Attacks. Gradient leakage attacks (Zhu et al., 2019; Zhao et al., 2020) exploit gradient matching techniques to recover clients’ data by aligning uploaded gradients with those from dummy data, enabling high-fidelity reconstruction. Subsequent research (Geiping et al., 2020; Yin et al., 2021; Wei et al., 2020) has improved the efficacy of these attacks by utilizing cosine similarity loss functions, adding regularization terms, and refining initialization methods, among others. Interestingly, optimized dummy images often resemble random noise, which hinders attack performance (Jeon et al., 2021). To address this, Jeon et al. (2021) and Yue et al. (2023) optimized generative models’ latent vectors to generate clearer images. Deng et al. (2021) found that combining $L_{1}$ and $L_{2}$ loss functions can improve text reconstruction. Balunovic et al. (2022) proposed a unified Bayesian framework for gradient leakage attacks, noting that larger models are more susceptible to privacy leakage. Besides, DMs and LLMs tend to be more susceptible to gradient leakage attacks because they directly model the training data. For example, Gupta et al. (2022) leveraged the generative capabilities of these models to generate candidate data and determined the optimal reconstruction by comparing the similarity between candidate gradients and the uploaded gradients. As model performance improves, so does the quality of reconstruction.
@@
Gradient Leakage Defenses. Encryption and perturbation are the primary techniques used to defend against gradient leakage attacks. The encryption methods (Zhang et al., 2020b; Wagh et al., 2020) prevent attackers from accessing individual client gradients, allowing only access to the plaintext of aggregated gradients, which forces them to reconstruct all clients’ data at once and complicates gradient matching problem. However, encryption methods can introduce a high computational overhead, limiting their usability in resource-constrained settings (Sun et al., 2021; Yue et al., 2023). Perturbation-based methods apply slight gradient modifications to confound attackers. Differential privacy (Abadi et al., 2016) injects random noise into gradients, while Top-K gradient sparsification (Zhu et al., 2019) retains only the most significant gradient elements. Gradient quantization (Yue et al., 2023) represents the gradients with lower-bit precision, reducing the amount of sensitive information. Selective pruning evaluates the private information within each gradient element and then implements gradient pruning (Sun et al., 2021). Another approach is to generate data that contains less private information, typically by optimizing metrics related to data privacy and model utility (Fan et al., 2022a).
@@
Remark. Gradient leakage attacks, initially designed for discriminative models, can also be applied to generative models, with DMs and LLMs being particularly vulnerable due to their ability to replicate training data patterns. As models become more proficient, the risk of disclosing training data also increases. In light of this, one possible mitigation involves commencing training with sensitive data and switching to less sensitive data; however, the impact of this strategy on model performance remains uncertain. For LLMs, RAG can be employed to store privacy-sensitive information locally, keeping it out of the training set for privacy protection. Furthermore, fine-tuning LLMs with adapters, rather than training from scratch, allows for a restricted sharing of gradients, reducing the risk of privacy breaches. Nonetheless, the effectiveness of the above several mitigations in counterbalancing the privacy risks introduced by models’ generation capabilities remains unclear.
@@ 3.2.2 Split Learning
In split learning, a model is divided into two sub-models: the bottom network on the client side and the top network on the server side. Clients process input data through the bottom network, sending intermediate results to the server for further computation using the top network. Privacy concerns arise when one party is untrustworthy.
Data Leakage. Pasquini et al. (2021) introduced Feature Space Hijacking Attack (FSHA), targeting the reconstruction of client data when the label party is untrustworthy. In FSHA, the adversarial label party trains a shadow network to mimic the outputs of client’s network, and then a decoder maps these outputs back to their corresponding data. By applying outputs from client’s bottom network to the trained decoder, client data can be revealed. Chen et al. (2024b) relaxed FSHA’s requirements, allowing the use of common public data instead of data similar to client data. Xu et al. (2024c) showed that replacing the decoder with a diffusion model improves reconstruction quality.
Existing defense measures, such as detection and perturbation, are found to be ineffective against FSHA (Pasquini et al., 2021). To address this, Li et al. (2022a) developed ResSFL, which trains a feature extractor resistant to inversion and initializes the bottom model with it. Maeng et al. (2024) theoretically analyzed the decoder’s reconstruction limits using Fisher information. Luo et al. (2023) employed a regularization term to reduce the correlation between the input data and intermediate activation values, alongside pruning sensitive parameters for defense.
Label Leakage. Li et al. (2022b) showed that backward gradient sign patterns from the top model can leak labels and suggested using gradient perturbations to mitigate. However, Xiao et al. (2021) argued against the effectiveness of this method and advocated for multiple activations and label mixing instead. Erdogan et al. (2021) and Kariyappa and Qureshi (2021) showed that labels could be inferred with gradient matching, and Fu et al. (2022) found that even a small amount of labeled samples is sufficient to fine-tune models, enabling direct label predictions for training data. Wan et al. (2023b) proposed using flipped labels to compute gradients to obfuscate attackers while training a private sub-model with true labels to compensate for performance losses.
@@ 3.3.2 LLMs
Association attacks. This type of attack involves designing a template, which is then filled with entity names to prompt the model to predict the corresponding sensitive information. As an example, Lehman et al. (2021) studied the model’s vulnerability to exposing sensitive medical information using a template: ”[NAME] symptoms of [masked].” Huang et al. (2022) examined the risk of leaking personal information, such as email addresses and phone numbers. Kim et al. (2023b) enhanced association attacks by employing likelihood ratio scores for more accurate predictions. Recent studies have also explored how the integration of external retrieval data affects privacy leakage in RAG. These works (Huang et al., 2023a; Zeng et al., 2024; Qi et al., 2024; Yu et al., 2023a) crafted specific templates to compel LLMs to output the retrieved data. The templates generally consist of prompts specifying the attacker-desired content and a command directing the model to present the retrieved content.
Defenses. Data deduplication (Carlini et al., 2021; Jagielski et al., 2023; Carlini et al., 2023b; Kandpal et al., 2022) and differential privacy (Carlini et al., 2021; Lukas et al., 2023) are both applicable to LLMs and present remarkable effectiveness in practice. Nevertheless, it is worth noting that differential privacy does come with the privacy-utility trade-off. For further protection, Lukas et al. (2023) suggested utilizing named entity recognition as a method to filter out sensitive information present in the training sets. This sensitive-information-filtering method works well in RAG (Huang et al., 2023a). Moreover, another approach to boost privacy protection in RAG is by blending public and private data in datastore and encoder training (Huang et al., 2023a).